OpenStudy (barrelracer011):

When 27 times x^2 times z all over negative 3 times x^2 z to the fourth power is completely simplified, the exponent on the variable z is _____

OpenStudy (anonymous):

Hold on gimme a sec

OpenStudy (barrelracer011):

Ok

OpenStudy (anonymous):

-9/z^5 I think

OpenStudy (barrelracer011):

So what would z be?

OpenStudy (anonymous):

Hmm not totally sure

OpenStudy (barrelracer011):

I need Z lol

OpenStudy (anonymous):

Hmm I know you can ask @texaschic101 to help

OpenStudy (johnweldon1993):

\[\large \frac{27x^2z}{-3x^2z^4}\] We just want the exponent on variable 'z' so focus on that \[\large \frac{z}{z^4} = z^{1 - 4} = z^{?}\]

OpenStudy (barrelracer011):

-3

OpenStudy (johnweldon1993):

Right...now I'm not sure what they mean by completely simplified... we can have \(\huge \frac{-9}{z^{3}}\) or we can have \(\large -9z^{-3}\)

OpenStudy (barrelracer011):

Im so confused

OpenStudy (johnweldon1993):

\[\large \frac{27\cancel{x^2}z}{-3\cancel{x^2}z^4}\] so all we really have is \[\large \frac{27z}{-3z^4}\] If we divide the 27/-3 we have -9...so now we have \[\large -9 \times \frac{z}{z^4}\]

OpenStudy (barrelracer011):

So it would be 4?

OpenStudy (johnweldon1993):

Now we just need to remember that \[\large \frac{n^a}{n^b} = n^{a - b}\] so \[\large \frac{z}{z^4} = z^{1 - 4} = z^{-3}\]

OpenStudy (johnweldon1993):

So altogether we have \(\large -9z^{-3}\) The exponent that goes with the 'z' would be -3 \[\large -9z^{\color \red{-3}}\]

OpenStudy (barrelracer011):

Yes but I put -3 last time and it was wrong

OpenStudy (johnweldon1993):

Well then maybe they want to keep this as \[\large \frac{\cancel{z^1}}{z^{\cancel{4} \rightarrow 3}} = \frac{1}{z^3}\]

OpenStudy (barrelracer011):

So it would be positive 3

OpenStudy (johnweldon1993):

Yeah if they don't want -3 then they didnt want that final simplification step...so yeah 3 would be correct then

OpenStudy (barrelracer011):

Ok thanks

OpenStudy (barrelracer011):

Yes it was positive 3

OpenStudy (johnweldon1993):

Alright that's what they wanted ....great! and anytime :)